# Thermoelectric coefficients of two-flavor quark matter from the Kubo formalism

### v1

@@ IV Numerical results
Before presenting our key results, we briefly review the generic features of the two-flavor NJL phase diagram and the scattering processes relevant for the quark spectral functions. For further details, see Refs. 1961PhRv..124..246N ; 1961PhRv..122..345N ; 1991PrPNP..27..195V ; 1992RvMP…64..649K ; 2005PhR…407..205B ; LKW15 ; Harutyunyan2017 .
@@ IV.1 Quark spectral functions
Figure 4 shows the phase diagram of the two-flavor NJL model on the $\mu$-$T$ plane within the approximations used. The shaded area separates the part of the phase diagram where the dispersive effects of the meson decay into on-shell quark-antiquark pair $\pi,\sigma\to q+\bar{q}$ are kinematically allowed. The inner boundary of the shaded area is the so-called Mott temperature $T_{\rm M}$ at which the pion mass equals twice the quark mass. In the case of exact chiral symmetry, i.e., $m_{0}=0$, the Mott temperature coincides with the critical temperature $T_{c}$ of the chiral phase transition. Above this temperature, we have $\langle\bar{\psi}\psi\rangle=0$ and $m=0$, signaling the restoration of chiral symmetry. The upper boundary of the diagram is the maximal temperature $T_{\rm max}$ above which mesonic modes do not exist anymore within the zero-momentum pole approximation of the ${\cal O}(N_{c})$ expansion, see Ref. Harutyunyan2017 for details. At $T=0$ this line ends at the maximal value of the chemical potential $\mu_{\rm max}=\Lambda$, where the meson masses are $m_{M}=2\Lambda$.
@@
The three components of the quark spectral function $A_{s}$, $A_{0}$ and $A_{v}$ can be expressed in terms of the relevant components of the quark self-energy according to the relations LW14 ; LKW15 ; Harutyunyan2017
@@
In this work, we will employ the quark spectral functions derived from one-meson exchange diagrams in the regime close to the chiral phase transition line. The imaginary parts of the on-shell quark and antiquark self-energies which arise from the meson decay into quark-antiquark pair were computed in Refs. LW14 ; LKW15 ; Harutyunyan2017
@@
Figure 5: Dependence of three Lorentz components of the quark spectral functions $a=-mA_{s}$ (solid lines), $b=-\varepsilon A_{0}$ (dash-dotted lines) and $c=-pA_{v}$ (dashed lines) on the quark momentum at $\mu=0$ for $T=0.22$ GeV (a) and $T=0.3$ GeV (b). These spectral functions are shown at three energies $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$, and $\varepsilon_{3}=0.5$ GeV, as indicated in the plot.
@@
The three Lorentz components of the spectral function obtained in Refs. LKW15 ; Harutyunyan2017 are shown in Fig. 5 as functions of the quark momentum for three values of the quark (off-shell) energy: $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$ and $\varepsilon_{3}=0.5$ GeV. The chemical potential is fixed at $\mu=0$; however, the general features of the spectral functions also remain the same for finite values of $\mu$. We see that the spectral functions are peaked around $p\simeq\varepsilon$, but the peaks are broadened with increasing temperature as more phase space is opened for the quark-meson scattering with increasing $T$.
@@ IV.2 Thermoelectric coefficients
To evaluate thermoelectric coefficients, we need the enthalpy per particle $h=\mu+Ts/n$, which within the $1/N_{c}$ (quasiparticle) approximation is given by the formulas 2005PhR…407..205B ; Harutyunyan2017
@@
The dependence of the conductivities $\sigma$ and $\tilde{\kappa}$ on $T$ and $\mu$ was plotted and discussed in detail in Ref. Harutyunyan2017 . These quantities were shown to decrease with increasing temperature for arbitrary values of the chemical potential, as a result of the broadening of the spectral functions. The same argument also holds for Eq. (73). In particular, they all diverge in the vicinity of the Mott temperature, because in the limit $T\to T_{M}$ the decay cross section of the mesons tends to zero, as a result of which the spectral functions transform into $\delta$-functions and lead to infrared divergence in Eqs. (72)–(III).
@@
Figure 6 (a) shows the temperature dependence of the modulus of thermopower Q-Q for constant values of the chemical potential. As $Q$ is given by the ratio of two correlation functions (73) and (72), their dependencies on the spectral functions, as well as the divergencies in the vicinity of the Mott temperature, are almost completely cancelled. As a result, $Q$ depends weakly on the microscopic interactions and has thermodynamic nature. We find that Q-Q increases almost linearly with increasing temperature. Approximating the integrands in formulas (73) and (72) as $pA_{v}\simeq\varepsilon A_{0}\gg mA_{s}$, we obtain the following simple approximation:
@@
where $\bar{\varepsilon}$ is the average characteristic quark energy. The second approximation is applicable in the case where $T\gtrsim\mu$, where $h\simeq 7\pi^{2}T^{2}/15\mu\geq\bar{\varepsilon}$ Harutyunyan2017 . In this regime, the absolute value of the thermopower increases linearly with temperature, in good agreement with the results shown in Fig. 6. From Eq. (90), we also see that $Q$ is negative at small to moderate values of $\mu$, but becomes positive at larger values of $\mu$, which are not considered here. It is worth noting that the increase of $Q$ with temperature in our model contrasts with the results of Refs. Dey2020 ; Singh2025 ; Shaikh2025 ; Rath2025 , where a perturbative QCD treatment within the Boltzmann equation framework was employed. However, we find very good agreement, both qualitative and quantitative, between our results for the thermopower and those of Ref. Abhishek2020 , which employed the same NJL model with quark–meson scattering above the Mott temperature, but used Boltzmann theory for weakly coupled quasiparticles to compute the transport coefficients.
@@
Nevertheless, their results for the electrical and thermal conductivities differ significantly from our findings discussed in Ref. Harutyunyan2017 . The agreement between the thermopower $Q$ in weakly and strongly interacting regimes, despite the strong discrepancy in conductivities, can be attributed to the fact that the thermopower is a ratio of two conductivities. As a result, the influence of the strength of microscopic interactions on $Q$ is relatively weak. We also note that the thermoelectric contribution to the thermal conductivity, $\tilde{\kappa}-\kappa=TQ^{2}\sigma$, which was neglected in Ref. Harutyunyan2017 , provides only a small correction of the order of 10–15%.
@@
where the derivative $({\partial\mu}/{\partial\beta})_{n}$ can be computed from Eq. (89Harutyunyan2017
@@
where $\tilde{n}^{\pm}\equiv 1-n^{\pm}$. Numerically, we find that the first term in Eq. (91) accounts for about one-third of the total $\rho$. As $Q\propto T$ for small chemical potentials, then the first term in Eq. (91) is almost equal to Q-Q, and we obtain a rough estimate $\rho\simeq-3Q$. Consequently, the Thomson effect becomes stronger at higher temperatures: for the same current and temperature gradient, more heat is released or absorbed. Note also that the Thomson coefficient is always positive in the regime of interest.
@@
The dependence of thermopower and the Thomson coefficient on the chemical potential for fixed values of the temperature is shown in Fig. 7. Both coefficients decrease with the chemical potential and tend to infinity in the limit $\mu\to 0$. The observed divergence at vanishing chemical potential is similar to the behavior of the thermal conductivity Harutyunyan2017 , and arises due to the terms containing $h$ in expressions (73) and (III). Within the domain of temperatures and chemical potentials of investigation, the minimum value of enthalpy per particle is nearly $h_{\rm min}\simeq 0.8$ GeV Harutyunyan2017 , which exceeds the cutoff parameter $\Lambda\simeq 0.65$ GeV, which serves as a characteristic energy scale for quarks. Thus, it can be concluded that the dominant terms in $Q$ and $\rho$ are those containing $h$. The physical reason for this divergence is the disappearance of the rest frame of the particles in the $n\to 0$ limit, relative to which the thermal current is defined 1985PhRvD..31…53D .
@@
Using the numerical results obtained above, we can make rough estimates of electric fields which will be generated in quark matter due to temperature gradients. A simple estimate gives $E\simeq T|Q|/l$, where $l$ is the characteristic average length scale of thermal gradients. Substituting Eq. (90), we obtain
@@
where in the second step we assumed a value $l\simeq 10$ fm typical for heavy-ion collisions. For typical values $0.2\leq T\leq 0.3$ GeV and $0.1\leq\mu\leq 0.2$ GeV, we find $1.2\cdot 10^{-2}\lesssim eE\lesssim 5.4\cdot 10^{-2}$ GeV2, or, in more relevant $m_{\pi}^{2}\simeq 2\cdot 10^{-2}$ GeV2 units
@@
This estimate is close to the values of electric fields induced by thermoelectric effects reported in Ref. Singh2024 , where the hydrodynamic equations for an expanding QGP fireball created in head-on heavy-ion collisions were solved for both Bjorken and Gubser flow scenarios. These estimated electric fields underscore the importance of including thermoelectric effects in the phenomenology of heavy-ion collisions.

### v2

@@ IV Numerical results
Before presenting our key results, we briefly review the generic features of the two-flavor NJL phase diagram and the scattering processes relevant for the quark spectral functions. For further details, see Refs. Nambu and Jona-Lasinio (1961b, a); Vogl and Weise (1991); Klevansky (1992); Buballa (2005); Lang et al. (2015); Harutyunyan et al. (2017).
@@ IV.1 Quark spectral functions
Figure 4 shows the phase diagram of the two-flavor NJL model on the $\mu$-$T$ plane within the approximations used. The shaded area separates the part of the phase diagram where the dispersive effects of the meson decay into on-shell quark-antiquark pair $\pi,\sigma\to q+\bar{q}$ are kinematically allowed. The inner boundary of the shaded area is the so-called Mott temperature $T_{\rm M}$ at which the pion mass equals twice the quark mass. In the case of exact chiral symmetry, i.e., $m_{0}=0$, the Mott temperature coincides with the critical temperature $T_{c}$ of the chiral phase transition. Above this temperature, we have $\langle\bar{\psi}\psi\rangle=0$ and $m=0$, signaling the restoration of chiral symmetry. The upper boundary of the diagram is the maximal temperature $T_{\rm max}$ above which mesonic modes do not exist anymore within the zero-momentum pole approximation of the ${\cal O}(N_{c})$ expansion, see Ref. Harutyunyan et al. (2017) for details. At $T=0$ this line ends at the maximal value of the chemical potential $\mu_{\rm max}=\Lambda$, where the meson masses are $m_{M}=2\Lambda$.
@@
The three components of the quark spectral function $A_{s}$, $A_{0}$, and $A_{v}$ can be expressed in terms of the relevant components of the quark self-energy according to the relations Lang and Weise (2014); Lang et al. (2015); Harutyunyan et al. (2017)
@@
In this work, we will employ the quark spectral functions derived from one-meson-exchange diagrams in the regime close to the chiral phase transition line. The imaginary parts of the on-shell quark and antiquark self-energies which arise from the meson decay into quark-antiquark pair were computed in Refs. Lang and Weise (2014); Lang et al. (2015); Harutyunyan et al. (2017)
@@
Figure 5: Dependence of three Lorentz components of the quark spectral functions $a=-mA_{s}$ (solid lines), $b=-\varepsilon A_{0}$ (dash-dotted lines), and $c=-pA_{v}$ (dashed lines) on the quark momentum at $\mu=0$ for $T=0.22$ GeV (a) and $T=0.3$ GeV (b). These spectral functions are shown at three energies $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$, and $\varepsilon_{3}=0.5$ GeV, as indicated in the plot.
@@
The three Lorentz components of the spectral function obtained in Refs. Lang et al. (2015); Harutyunyan et al. (2017) are shown in Fig. 5 as functions of the quark momentum for three values of the quark (off-shell) energy: $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$, and $\varepsilon_{3}=0.5$ GeV. The chemical potential is fixed at $\mu=0$; however, the general features of the spectral functions also remain the same for finite values of $\mu$. We see that the spectral functions are peaked around $p\simeq\varepsilon$, but the peaks are broadened with increasing temperature as more phase space is opened for the quark-meson scattering with increasing $T$.
Table 1: Comparison of thermopower $Q$ obtained in different theoretical approaches. The quoted values correspond to typical magnitudes in the indicated temperature and chemical-potential ranges. For $Q(T)$ and $Q(\mu)$, “Increasing” (respectively, “Decreasing”) indicates that the thermopower grows (respectively, decreases) with temperature $T$ or chemical potential $\mu$. Center dots $\dots$ indicate that no clear trend is reported in the reference.
Ref. Model / approach d.o.f. $T$ range (GeV) $\mu$ range (GeV) $Q$ $Q(T)$ $Q(\mu)$
This work NJL/Kubo 2ff quarks + mesons 0.170.170.300.30 0.050.050.200.20 (-) 554545 Increasing Decreasing
Dey and Patra (2020) pQCD / RTA 3ff quarks + gluons 0.170.170.400.40 0.030.030.050.05 (+) 0.010.010.080.08 Decreasing Increasing
Kurian (2021) pQCD / RTA 3ff quarks + gluons 0.260.260.400.40 0.080.080.100.10 (-) 553535 Decreasing Increasing
Rath and Neill (2025) pQCD/ RTA 3ff quarks + gluons 0.170.170.640.64 0.040.040.080.08 (+) 0.010.010.100.10 Decreasing Increasing
Shaikh et al. (2025) pQCD / RTA 3ff quarks + gluons 0.170.170.650.65 0.040.040.080.08 (+) 0.010.011616 Decreasing Increasing
Khan et al. (2024) pQCD / RTA 3ff quarks + gluons 0.170.170.500.50 $0.04$ (+) 0.350.3511 Decreasing Increasing
Singh and Sahoo (2025) pQCD / RTA 2ff quarks + mesons 0.16–0.32 0.10–0.60 (-) 0.10.155 Increasing Increasing
Abhishek et al. (2022) NJL / RTA 2ff quarks + mesons 0.10–0.26 0.10–0.20 (-) 222020 Decreasing Increasing
Zhang et al. (2022) NJL / RTA 2ff quarks + mesons 0.10–0.40 0.10 (+) 7 Const. $\dots$
@@ IV.2 Thermoelectric coefficients
To evaluate thermoelectric coefficients, we need the enthalpy per particle $h=\mu+Ts/n$, which within the $1/N_{c}$ (quasiparticle) approximation is given by the formulas Buballa (2005); Harutyunyan et al. (2017)
@@ IV.2.1 Thermopower
The dependence of the conductivities $\sigma$ and $\tilde{\kappa}$ on $T$ and $\mu$ was plotted and discussed in detail in Ref. Harutyunyan et al. (2017). These quantities were shown to decrease with increasing temperature for arbitrary values of the chemical potential, as a result of the broadening of the spectral functions. The same argument also holds for Eq. (73). In particular, they all diverge in the vicinity of the Mott temperature, because in the limit $T\to T_{M}$ the decay cross section of the mesons tends to zero, as a result of which the spectral functions transform into $\delta$-functions and lead to infrared divergence in Eqs. (72)–(III).
@@
Figure 6 (a) shows the temperature dependence of the modulus of thermopower Q-Q for constant values of the chemical potential. As $Q$ is given by the ratio of two correlation functions (73) and (72), their dependencies on the spectral functions, as well as the divergencies in the vicinity of the Mott temperature, are almost completely canceled. As a result, $Q$ depends weakly on the microscopic interactions and has thermodynamic nature. We find that Q-Q increases almost linearly with increasing temperature. Approximating the integrands in formulas (73) and (72) as $pA_{v}\simeq\varepsilon A_{0}\gg mA_{s}$, we obtain the following simple approximation:
@@
where $\bar{\varepsilon}$ is the average characteristic quark energy. The second approximation is applicable in the case where $T\gtrsim\mu$, where $h\simeq 7\pi^{2}T^{2}/15\mu\geq\bar{\varepsilon}$ Harutyunyan et al. (2017). In this regime, the absolute value of the thermopower increases linearly with temperature, in good agreement with the results shown in Fig. 6. From Eq. (90), we also see that $Q$ is negative at small to moderate values of $\mu$, but becomes positive at larger values of $\mu$, which are not considered here.
Table 1 provides a comparison of the general trends and quantitative comparison of the thermopower obtained in different theoretical approaches Dey and Patra (2020); Kurian (2021); Rath and Neill (2025); Shaikh et al. (2025); Khan et al. (2024); Singh and Sahoo (2025); Abhishek et al. (2022); Zhang et al. (2022). Our results, computed within the NJL model above the Mott temperature, show relatively large negative values of $Q$ in the range (5)(-5)(45)(-45), with a clear trend of increasing magnitude with temperature and decreasing magnitude with chemical potential. This behavior contrasts with the perturbative quantum chromodynamics (pQCD) combined with Boltzmann-equation-based relaxation-time approximation (RTA) results reported in Refs. Dey and Patra (2020); Singh and Sahoo (2025); Shaikh et al. (2025); Khan et al. (2024); Kurian (2021), where the thermopower is generally smaller in magnitude, often positive, and exhibits the opposite qualitative trend with respect to $\mu$, i.e., $Q$ increases with chemical potential. In particular, for chemical potentials $\mu\leq 100$ MeV, our results for the absolute value of the Seebeck coefficient exceed those reported in Refs. Dey and Patra (2020); Rath and Neill (2025); Shaikh et al. (2025); Khan et al. (2024), and Singh and Sahoo (2025) by approximately 4, 3, 1, 2, and 2 orders of magnitude, respectively. Similarly, the Seebeck coefficients obtained in other NJL-based works Abhishek et al. (2022); Zhang et al. (2022) show trends comparable to ours, confirming the robustness of the RTA-NJL approach in the high-temperature, quark-meson phase of the model. The differences in the absolute value of $Q$ with pQCD/RTA calculations can be attributed to the inclusion of gluon degrees of freedom, the use of energy-dependent relaxation times, and the different microscopic scattering mechanisms underlying these approaches. Overall, our results highlight that the NJL framework predicts substantially larger thermoelectric response in the quark-meson plasma than pQCD-based models. It is worth noting that the increase of $Q$ with temperature observed in our model contrasts with the results of Refs. Dey and Patra (2020); Rath and Neill (2025); Singh and Sahoo (2025); Shaikh et al. (2025); Khan et al. (2024) based on pQCD/RTA approach. The difference arises from the fact that the quasiparticle widths are parametrically small and scattering is dominated by leading-order perturbative processes, which tend to suppress thermoelectric response. In contrast, our approach incorporates nonperturbative effects through finite spectral widths of the quark propagators, leading to enhanced interaction rates and, consequently, larger values of the Seebeck coefficient.
Turning to the comparison with the NJL/RTA approaches, we note that the qualitative dependence of the thermopower found in this work on temperature and chemical potential is similar to that reported in Ref. Kurian (2021); however, the absolute magnitude of our results is larger by about a factor of 4, which can be traced to the different treatment of quasiparticle damping and medium effects.
@@
Furthermore, we find very good agreement, both qualitative and quantitative, between our results for the thermopower and those of Ref. Abhishek et al. (2022), which employed the same NJL model with quark–meson scattering above the Mott temperature, but used Boltzmann theory for weakly coupled quasiparticles to compute the transport coefficients. Nevertheless, the results of Abhishek et al. (2022) for the electrical and thermal conductivities differ significantly from our findings discussed in Ref. Harutyunyan et al. (2017). The agreement between the thermopower $Q$ in weakly and strongly interacting regimes, despite the strong discrepancy in conductivities, can be attributed to the fact that the thermopower is a ratio of two conductivities. As a result, the influence of the strength of microscopic interactions on $Q$ is relatively weak. We also note that the thermoelectric contribution to the thermal conductivity, $\tilde{\kappa}-\kappa=TQ^{2}\sigma$, which was neglected in Ref. Harutyunyan et al. (2017), provides only a small correction of the order of 10%–15%.
IV.2.2 The Thomson coefficient
@@
where the derivative $({\partial\mu}/{\partial\beta})_{n}$ can be computed from Eq. (89Harutyunyan et al. (2017)
@@
where $\tilde{n}^{\pm}\equiv 1-n^{\pm}$. Numerically, we find that the first term in Eq. (91) accounts for about one-third of the total $\rho$. As $Q\propto T$ for small chemical potentials, then the first term in Eq. (91) is almost equal to Q-Q, and we obtain a rough estimate $\rho\simeq-3Q$. Consequently, the Thomson effect becomes stronger at higher temperatures: for the same current and temperature gradient, more heat is released or absorbed. Note also that the Thomson coefficient is always positive in the regime of interest. This result contrasts with that of Ref. Singh and Sahoo (2025), where the Thomson coefficient may be either positive or negative, depending on the value of the chemical potential. Moreover, the magnitude of the Thomson coefficient obtained in our work, $|\rho|$, exceeds the values reported in Ref. Singh and Sahoo (2025) by 1–2 to orders of magnitude.
@@
The dependence of thermopower and the Thomson coefficient on the chemical potential for fixed values of the temperature is shown in Fig. 7. Both coefficients decrease with the chemical potential and tend to infinity in the limit $\mu\to 0$. The observed divergence at vanishing chemical potential is similar to the behavior of the thermal conductivity Harutyunyan et al. (2017) and arises due to the terms containing $h$ in expressions (73) and (III). Within the domain of temperatures and chemical potentials of investigation, the minimum value of enthalpy per particle is nearly $h_{\rm min}\simeq 0.8$ GeV Harutyunyan et al. (2017), which exceeds the cutoff parameter $\Lambda\simeq 0.65$ GeV, which serves as a characteristic energy scale for quarks. Thus, it can be concluded that the dominant terms in $Q$ and $\rho$ are those containing $h$. The physical reason for this divergence is the disappearance of the rest frame of the particles in the $n\to 0$ limit, relative to which the thermal current is defined Danielewicz and Gyulassy (1985).
Figure 8: The Lorenz ratio $L=\kappa/\sigma T$ (a) and the ratio $L/Q^{2}$ (b) as functions of the temperature at various values of the chemical potential indicated on the plots. The vertical lines show the Mott temperature at the given value of $\mu$.
Figure 9: The Lorenz ratio $L=\kappa/\sigma T$(a) and the ratio $L/Q^{2}$ (b) as functions of the chemical potential at various values of the temperature indicated on the plots. The vertical line shows the value of the chemical potential on the Mott line for temperature $T=0.2$ GeV.
IV.2.3 The Lorenz ratio
As shown earlier in Ref. Harutyunyan et al. (2017), the Wiedemann–Franz law for the Lorenz ratio $L=\kappa/(\sigma T)$ does not hold in hot relativistic quark matter in the regime considered in this work. This violation originates from the divergence of the thermal current at small chemical potentials. A simple estimate of the Lorenz ratio was also provided in Ref. Harutyunyan et al. (2017)
$L\simeq\frac{N_{f}}{{\rm Tr}\hat{q}^{2}}\frac{h^{2}}{T^{2}}\simeq\frac{98\pi^{4}}{125e^{2}}\frac{T^{2}}{\mu^{2}}.$ (93)
Although this estimate was derived for the regime $\mu\ll T$, the exact numerical results suggest that it is valid also for larger domain of chemical potentials $\mu\lesssim T$. This can be seen from Figs. 8(a) and 9(a), which show the exact Lorenz ratio as a function of temperature and chemical potential, respectively.
We recall that for classical nonrelativistic electron-ion plasma the Lorenz ratio has the constant value $L_{\rm WF}=3/2e^{2}$, according to the Wiedemann-Franz law. However, in a multicomponent systems such as QGP, the Lorenz ratio should be modified by substitution $e^{2}\to{\rm Tr}\hat{q}^{2}/N_{f}$ to account for different vales of electric charges of different quark species in the plasma Rai et al. (2025). Thus, the classical Lorenz ratio for two-flavor quark matter would be $L_{\rm WF}=27/5e^{2}\simeq 59$ (recall that in the units we use $e^{2}=4\pi\alpha$). From Figs. 8 and 9 we see that at $\mu\ll T$, the Lorenz ratio $L\gg L_{\rm WF}$, as found also in Rai et al. (2025). However, the Wiedemann-Franz law may be restored in the regime of degenerate quark matter with $\mu\gg T$, as mentioned in Refs. Lucas and Chung Fong (2018); Rai et al. (2025) (in this regime for two-flavor quark matter one obtains $L_{\rm WF}=6\pi^{2}/5e^{2}\simeq 129$, instead of its value $L_{\rm WF}=\pi^{2}/3e^{2}$ for ordinary metals).
Comparing Eqs. (90) and (93), we find a simple approximate relation $L/Q^{2}\simeq 10$. Figures 8(b) and 9(b) show the exact value of $L/Q^{2}$ as a function of temperature and chemical potential, respectively. We see that, in the regime where $\mu\ll T$, this ratio tends to a constant value $L/Q^{2}\to 9$, and its variations are small in the whole range of temperatures and chemical potentials considered here (note that in degenerate gas $L/Q^{2}\to\infty$ as $T\to 0$, because $Q\propto T$ in that regime Ziman (1979)).
IV.2.4 Generated thermoelectric fields
@@
Using the numerical results obtained above, we can make rough estimates of electric fields which will be generated in quark matter due to temperature gradients. A simple estimate gives $E\simeq T|Q|/l$, where $l$ is the characteristic average length scale of thermal gradients. Substituting here Eq. (90), we obtain
@@
where in the second step we assumed a value $l\simeq 10$ fm typical for heavy-ion collisions. For typical values $0.2\leq T\leq 0.3$ and $0.1\leq\mu\leq 0.2$ GeV, we find $1.2\cdot 10^{-2}\lesssim eE\lesssim 5.4\cdot 10^{-2}$ GeV2, or, in more relevant $m_{\pi}^{2}\simeq 2\cdot 10^{-2}$ GeV2 units
@@
This estimate is close to the values of electric fields induced by thermoelectric effects reported in Ref. Singh et al. (2024), where the hydrodynamic equations for an expanding QGP fireball created in head-on heavy-ion collisions were solved for both Bjorken and Gubser flow scenarios. These estimated electric fields underscore the importance of including thermoelectric effects in the phenomenology of heavy-ion collisions.